L-1(s2 - 4 / ( s+2) 2)


1
Expert's answer
2021-06-22T06:15:36-0400

Using the table of originals and images, we get:

"{L^{ - 1}}\left( {\frac{1}{{{{(s + 2)}^2}}}} \right) = t{e^{ - 2t}}"

By the image differentiation theorem

"{L^{ - 1}}\left( {{s^2}} \right) = \delta ''\left( t \right)"

Then

"{L^{ - 1}}\left( {{s^2} - \frac{4}{{{{(s + 2)}^2}}}} \right) = {L^{ - 1}}\left( {{s^2}} \right) - 4{L^{ - 1}}\left( {\frac{1}{{{{(s + 2)}^2}}}} \right) = \delta ''\left( t \right) - 4t{e^{ - 2t}}"

Answer: "\delta ''\left( t \right) - 4t{e^{ - 2t}}" , "\delta \left( t \right)" is Dirac delta function